# A Platform for Evanescently Trapping 87Rb Using Silicon Nitride Strip Waveguides Buried in Silica

### v1

@@ Abstract
Cold-atom systems have emerged as a highly promising avenue for quantum-enhanced position, navigation, and timing applications. However, their wider adoption is currently hampered in part by the large footprint of the systems. In leveraging the miniaturisation possible through photonic integrated circuits, cold-atom sensors would be able to reach much wider commercial adoption. In this paper, we introduce a platform for evanescently trapping 87Rb using strip silicon nitride waveguides buried in silica using red- and blue-detuned fundamental and higher-order modes, providing a three-dimensional adjustable trap for BEC-based, chip-scale work in quantum science and technologies.
@@ I INTRODUCTION
Ultracold atoms represent the future of precision measurement of time and acceleration. Since the first magneto-optical trap (MOT) [raab_trapping_1987] and the first experimental observation of a Bose-Einstein condensate (BEC) [anderson_observation_1995], research on engineering applications of ultracold matter has expanded significantly. Cesium fountain clocks remain the basis for the SI second [bipm_international_2019], and optical lattice clocks have achieved fractional uncertainties as low as $10^{-18}$ [mcgrew_atomic_2018]. Global navigation satellite systems rely on microwave transitions in rubidium and cesium atoms for timing and positioning and underpin modern navigation, while atom interferometry has enabled high-precision inertial sensing [kasevich_atomic_1991, dickerson_multiaxis_2013, weidner_experimental_2018, rosi_precision_2014] for gravimetry and other emerging applications.
@@ II.1 Wire Trapping
Table 1: Trap parameters for the gMOT, evaporative, and PIC trap to be transferred using the wire trap system. The typical trapping height from the chip is denoted by $z_{\textrm{trap}}$; $\Delta T=\mu_{B}\big[U(\textbf{r}_{\textrm{trap}})-U(\textbf{r}=\infty)\big]$ is the difference in the potential at the trap and far from the trap; and $\omega$ is the trap’s frequency.
Although numerical methods are typically employed to optimise dimple traps in experimental implementations, useful intuition can be gained from the analytic case of a single wire producing a 2D trap. This allows us to estimate the required magnitudes of current $I$ and bias field $B_{\textrm{ext}}$ to achieve a desired trap height $z_{\textrm{trap}}$, depth $\Delta T$, and trap frequency $\omega$.
@@
$\omega^{2}=\frac{\mu_{B}B_{\textrm{ext}}}{mz^{2}_{\textrm{trap}}}.$ (6)
@@
II.2 Transferring Between Traps
@@ III.1 Design of Vertical Trapping Potentials
Equation (14) is a transcendental equation that cannot be solved analytically for $z$, but numerically we can calculate what values of $U_{r,b}$ are required to produce a trap of depth $\Delta T_{\textrm{trap}}$ and a barrier $\Delta T_{\textrm{barrier}}$ (Table 2). Figure 4(a) shows the shape of this potential along the centre of the waveguide $y=0$, and a stable trap of $\Delta T=1$ µK is formed at $z=205$ nm. Figure 5 describes the mode powers incident down the trapping waveguide and their effect on the barrier/trap’s potential height/depth. An acceptability criteria of $0.8$ µK $\Delta T_{\textrm{trap}}<1.2$ µK and 88 µK $\Delta T_{\textrm{barrier}}<20$ µK is chosen and highlighted to illustrate this system’s tunability and robustness to fluctuations in laser power.
@@ III.3 Trap Properties
Table 3: Numerical calculations of the trap frequencies in all dimensions, positions, and depth at $\textbf{r}_{\textrm{lattice}}$ and $\textbf{r}_{\textrm{saddle}}$. These locations are independently calculated, as $z_{\textrm{lattice}}\neq z_{\textrm{saddle}}$, with $z_{\textrm{saddle}}=380$ nm. A negative value of $f_{x}(\textbf{r}_{\textrm{saddle}})$ illustrates the saddle potential’s steepness. Agreement within 10% was found between numerical and analytical expressions.
@@ III.4 Component Design
PIC mode converters are also required to construct the trap. The total blue mode intensity (TE0 + TE1) is 13.4mW, with about 78% and 22% needed in each mode respectively. Asymmetrical waveguides and directional coupler mode converters are able to achieve 95% mode-conversion efficiency [gallacher_silicon_2022, sulway_novel_2022]; meaning this fraction is achievable to engineer, as full mode conversion is not required.
@@
FIG. 7: (a) Coupling lengths $L_{c}$ for red TE0, blue TE0, and blue TE1 modes against DC separation $s_{DC}$. Fractional comparisons to the blue TE0 mode are also plotted, where advantageously $L_{C}(\lambda=720\text{nm, TE0})\approx 2L_{C}(\lambda=\text{850nm, TE0})$, but regrettably $L_{C}(\lambda=\text{720nm, TE0})\approx 4L_{C}(\lambda=\text{720nm, TE1})$. (b) An interaction length $\Delta x=695$nm is chosen, and the effect on how many $L_{c}$ each mode experiences is plotted. An $s_{DC}=100$nm is chosen, which means each mode goes through $(1.1,2.2,4.8)L_{c}$ respectively. This yields a total loss $L=0.51\text{dB}$.
@@ IV.1 Stray Charges
FIG. 8: Histogram showing the effective temperature shift induced in the vertical direction by stray charges over $10^{4}$ instances. Charges are randomly distributed across the surface in lab conditions (charge density of $\sigma=10^{-6}$C/m2). This is considered at the barrier (a,b); trap (c,d); over a large range (a,c); and the worst 10% of instances (b,d).
@@ IV.2 Chiral Effects (or lack thereof)
A consideration frequently required in atom traps using strong light confinement is that of chiral quantum optics [lodahl_chiral_2017]. The emergent effect appears with waveguides of sub-$\lambda$ geometries ($w_{\textrm{wg}}<\lambda/n_{eff}$), where the local polarisation can be dependent to the propagation direction of the light, hence producing chiral fields. This can lead to inhomogeneous Zeeman broadening which arise from the vector light shifts which in turn make the trap state-dependent and spatially inhomogeneous. In fiber-based applications this is typically avoided by using counter-propagating modes and considering the light-shift Hamiltonian with respect to the polarisability tensor $\alpha^{(2)}(\omega)$ [goban_demonstration_2012, pache_magic-wavelength_2025]. In the system presented, modes are polarised along the TE0 direction, where $w_{\textrm{wg}}$ is significantly greater than $\lambda$; leading to minimal chiral effects to begin with and a trap that is proportionally much more localised (minimising spatial inhomogeneity). Thus, only trapping and transport modes are required within the PIC, and no modes are needed to remove the vector light shift; greatly simplifying the experimental complexity of the system.
@@ V.2 Inertial Sensing
Using a single, one-dimensional waveguide, the methods considered here will largely be limited to accelerometry. However, given the compact size of the device, one can imagine an array of waveguides in orthogonal directions, which provides both options for relatively high-bandwidth atom interferometry (by running multiple PIC-based interferometers loaded from a single BEC) and multi-axis inertial sensing of accelerations and rotations. Given that chip-scale BECs have been produced at rates of $2\text{\,}\mathrm{Hz}$ [klempt], such serial operation of a set of independent interferometers can, in principle, reach repetition rates in the tens to hundreds of interferometer cycles per second.

### v2

# A Platform for Evanescently Trapping $^{\textrm{87}}$Rb Using Silicon Nitride Strip Waveguides Buried in Silica

@@ Abstract
Cold-atom systems have emerged as a highly promising avenue for quantum-enhanced position, navigation, and timing applications. However, their wider adoption is currently hampered in part by the large footprint of the systems. In leveraging the miniaturisation possible through photonic integrated circuits, cold-atom sensors would be able to reach much wider commercial adoption. In this paper, we introduce a platform for evanescently trapping $^{\textrm{87}}$Rb using strip silicon nitride waveguides buried in silica using red- and blue-detuned fundamental and higher-order modes, providing a three-dimensional adjustable trap for BEC-based, chip-scale work in quantum science and technologies.
Ultracold atoms, Photonic integrated circuits, Inertial sensing, Atomic clocks
@@ I Introduction
Ultracold atoms represent the future of precision measurement of time and acceleration. Since the first magneto-optical trap (MOT) [raab_trapping_1987] and the first experimental observation of a Bose-Einstein condensate (BEC) [anderson_observation_1995], research on engineering applications of ultracold matter has expanded significantly. Caesium fountain clocks remain the basis for the SI second [bipm_international_2019], and optical lattice clocks have achieved fractional uncertainties as low as $10^{-18}$ [mcgrew_atomic_2018]. Global navigation satellite systems rely on microwave transitions in rubidium and caesium atoms for timing and positioning and underpin modern navigation, while atom interferometry has enabled high-precision inertial sensing [kasevich_atomic_1991, dickerson_multiaxis_2013, weidner_experimental_2018, rosi_precision_2014] for gravimetry and other emerging applications.
@@ II.2 Wire Trapping
Table 1: Trap parameters for the gMOT, evaporative, and PIC trap to be transferred using the wire trap system. The typical trapping height from the chip is denoted by $z_{\textrm{trap}}$; $\Delta T=\mu_{B}\big[U(\textbf{r}_{\textrm{trap}})-U(\textbf{r}=\infty)\big]$ is the difference in the potential at the trap and far from the trap; and $\omega_{\textrm{trap}}$ is the trap’s frequency.
Although numerical methods are typically employed to optimise dimple traps in experimental implementations, useful intuition can be gained from the analytic case of a single wire producing a 2D trap. This allows us to estimate the required magnitudes of current $I$ and bias field $B_{\textrm{ext}}$ to achieve a desired trap height $z_{\textrm{trap}}$, depth $\Delta T$, and trap frequency $\omega_{\textrm{trap}}$.
@@
$\omega_{\textrm{trap}}^{2}=\frac{\mu_{B}B_{\textrm{ext}}}{mz^{2}_{\textrm{trap}}}.$ (6)
@@
II.3 Transferring Between Traps
@@ III.1 Design of Vertical Trapping Potentials
Equation (14) is a transcendental equation that cannot be solved analytically for $z$, but numerically we can calculate what values of $U_{r,b}$ are required to produce a trap of depth $\Delta T_{\textrm{trap}}$ and a barrier $\Delta T_{\textrm{barrier}}$ (Table 2). Figure 4(a) shows the shape of this potential along the centre of the waveguide $y=0$, and a stable trap of $\Delta T=1$ µK is formed at $z=205$ nm. Figure 5 describes the mode powers incident down the trapping waveguide and their effect on the barrier/trap’s potential height/depth. We adopt acceptability criteria of trap depth $7(\pm 1)E_{\textrm{rec}}$ and $30(\pm 3)E_{\textrm{rec}}$, together with a barrier potential of 88 µK $\Delta T_{\textrm{barrier}}<20$ µK. These regions are highlighted to illustrate this system’s tunability and robustness to fluctuations in laser power.
@@ III.3 Trap Properties
Table 3: Numerical calculations of the trap frequencies in all dimensions, positions, and depth at $\textbf{r}_{\textrm{lattice}}$ and $\textbf{r}_{\textrm{saddle}}$. These locations are independently calculated, as $z_{\textrm{lattice}}\neq z_{\textrm{saddle}}$, with $z_{\textrm{saddle}}=380$ nm. A negative value of $f_{x}(\textbf{r}_{\textrm{saddle}})$ illustrates the saddle potential’s steepness.
@@ III.4 Component Design
PIC mode converters are also required to construct the trap. The total blue mode intensity (TE0 + TE1) is 13.4mW, with about 78% and 22% needed in each mode respectively. Asymmetrical single-waveguides and directional coupler mode converters are able to achieve 95% mode-conversion efficiency [gallacher_silicon_2022, sulway_novel_2022]; meaning this fraction is achievable to engineer, as full mode conversion is not required.
@@
FIG. 7: (a) Coupling lengths $L_{c}$ for red TE0, blue TE0, and blue TE1 modes against DC separation $s_{DC}$. Fractional comparisons to the blue TE0 mode are also plotted, where advantageously $L_{C}(\lambda=720\text{nm, TE0})\approx 2L_{C}(\lambda=\text{850nm, TE0})$, but regrettably $L_{C}(\lambda=\text{720nm, TE0})\approx 4L_{C}(\lambda=\text{720nm, TE1})$. (b) An interaction length $\Delta x=695$nm is chosen, and the effect on how many $L_{c}$ each mode experiences is plotted. An $s_{DC}=100$nm is chosen, which means each mode goes through $\{1.1,2.2,4.8\}\,L_{c}$ respectively. This yields a total loss $L=0.51\text{dB}$.
@@ IV.1 Stray Charges
FIG. 8: Histogram showing the effective temperature shift induced in the vertical direction by stray charges over $10^{5}$ instances. Charges are randomly distributed across the surface in lab conditions (charge density of $\sigma=10^{-6}$C/m2). This is considered at the barrier (a,b); trap (c,d); over a large range (a,c); and the worst 10% of instances (b,d).
@@ IV.2 Chiral Effects (or lack thereof)
A consideration frequently required in atom traps using strong light confinement is that of chiral quantum optics [lodahl_chiral_2017]. The emergent effect appears with waveguides of sub-$\lambda$ geometries ($w_{\textrm{wg}}<\lambda/n_{\textrm{eff}}$), where the local polarisation can be dependent to the propagation direction of the light, hence producing chiral fields. This can lead to inhomogeneous Zeeman broadening which arise from the vector light shifts which in turn make the trap state-dependent and spatially inhomogeneous. In fiber-based applications this is typically avoided by using counter-propagating modes and considering the light-shift Hamiltonian with respect to the polarisability tensor $\alpha^{(2)}(\omega)$ [goban_demonstration_2012, pache_magic-wavelength_2025]. In the system presented, modes are polarised along the TE0 direction, where $w_{\textrm{wg}}$ is significantly greater than $\lambda$; leading to minimal chiral effects to begin with and a trap that is proportionally much more localised (minimising spatial inhomogeneity). Thus, only trapping and transport modes are required within the PIC, and no modes are needed to remove the vector light shift; greatly simplifying the experimental complexity of the system.
@@ V.2 Inertial Sensing
Using a single, one-dimensional waveguide, the methods considered here will largely be limited to accelerometry, although similar proposals for rotation sensing have been recently developed [zhouSagnacTractorAtom2025]. However, given the compact size of the device, one can imagine an array of waveguides in orthogonal directions, which provides both options for relatively high-bandwidth atom interferometry (by running multiple PIC-based interferometers loaded from a single BEC) and multi-axis inertial sensing of accelerations and rotations. Given that chip-scale BECs have been produced at rates of $2\text{\,}\mathrm{Hz}$ [klempt], such serial operation of a set of independent interferometers can, in principle, reach repetition rates in the tens to hundreds of interferometer cycles per second.
@@ VI Conclusion
During the preparation of this work, the authors became aware of work by Zhou et al. [zhouSagnacTractorAtom2025] that describes a similar method for atom trapping on a PIC. Their method focuses on inertial sensing using atoms trapped in ring waveguides with somewhat different loading conditions, and this work focuses on accelerometry and timekeeping, with the final goal of producing a self-contained ultracold atom sensor.